OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24> Please help I have no idea how to do this....

OpenStudy (michele_laino):

what is the scalar product of your vectors?

OpenStudy (anonymous):

parrell is when it would be -2/6 equals 24/8 orthogonal is when u x v is equal to zero

OpenStudy (anonymous):

8x6=48 and 24x-2=-48 so 48+-48 =0 so they are orthogonal

OpenStudy (anonymous):

okay! thank you so whats @lex245 whats an example of neither??

OpenStudy (anonymous):

@Michele_Laino I'm not sure what that is...

OpenStudy (anonymous):

parallel would be like <5,10> and <4,8> because 10/5 = 2 and 8/4=2

OpenStudy (anonymous):

orthogonal is when the dot product equals zero

OpenStudy (anonymous):

just like the question you originally asked

OpenStudy (anonymous):

okay, i get that now. :) thank you!!!

OpenStudy (anonymous):

no problem

OpenStudy (michele_laino):

by definition, the sclar product of your vectors, is the subsequent quantity: 6*8+(-2)*24=...?

OpenStudy (michele_laino):

scalar*

OpenStudy (anonymous):

@Michele_Laino can you solve for all answers \[[0, 2\Pi)\]\[3\tan ^{4}\alpha = 1 +\sec ^{2}\alpha]

OpenStudy (michele_laino):

please, apply this identity: \[{\left( {\tan \alpha } \right)^4} = {\left( {\frac{{\cos \alpha }}{{\sin \alpha }}} \right)^4} = \frac{{{{\left( {1 - {{\left( {\sin \alpha } \right)}^2}} \right)}^2}}}{{{{\left( {\sin \alpha } \right)}^2}}}\]